OpenStudy (anonymous):

Solve the equation 2/5 (x − 4) = 2x for x

OpenStudy (kinggeorge):

Is that \[\frac{2}{5}(x-4)=2x\]Or is it\[\frac{2}{5(x-4)}=2x\]

OpenStudy (anonymous):

IDK BRRO

OpenStudy (anonymous):

the first one

OpenStudy (kinggeorge):

I think we should probably assume it's the first one. Unless you're in the middle of a section on quadratic equations.

OpenStudy (kinggeorge):

First up, multiply both sides by 5\[\frac{2}{5}(x-4)=2x\implies2(x-4)=10x\]Distribute, and we get \[2x-8=10x\]Can you finish it from here?

OpenStudy (anonymous):

yeahh thnx medal 4 you!!!

OpenStudy (kinggeorge):

You're welcome. And thanks :)

OpenStudy (anonymous):

no prob dude!!